# First Measurement of Solar 8B Neutrinos via Coherent Elastic Neutrino-Nucleus Scattering with XENONnT

### v1

@@
First Measurement of Solar 8B Neutrinos via Coherent Elastic Neutrino-Nucleus Scattering with XENONnT
@@ Abstract
We present the first measurement of nuclear recoils from solar 8B neutrinos via coherent elastic neutrino-nucleus scattering with the XENONnT dark matter experiment. The central detector of XENONnT is a low-background, two-phase time projection chamber with a 5.9 t sensitive liquid xenon target. A blind analysis with an exposure of 3.51 ty resulted in 37 observed events above 0.5 keV, with () events expected from backgrounds. The background-only hypothesis is rejected with a statistical significance of 2.73 . The measured 8B solar neutrino flux of is consistent with results from dedicated solar neutrino experiments. The measured neutrino flux-weighted CENS cross-section on Xe of is consistent with the Standard Model prediction. This is the first direct measurement of nuclear recoils from solar neutrinos with a dark matter detector.
@@
Coherent elastic neutrino-nucleus scattering (CENS) is a Standard Model (SM) process with low momentum transfer, which allows neutrinos to scatter coherently with nuclei [1, 2]. This process has only recently been observed using an intense, pulsed spallation neutron source (SNS) [3, 4]. The detection of CENS events from solar neutrinos is more challenging due to the lower flux [5] and energy, and the lack of timing information. Therefore, it requires minimal backgrounds and maximizing the sensitive region of interest (ROI) with a low energy threshold. Liquid xenon (LXe) detectors searching for dark matter (DM) particles fulfill these requirements, but have not been able to reach the required sensitivity until now [6, 7]. Solar 8B neutrinos are expected to contribute the largest detectable number of coherent neutrino-xenon scattering events, albeit at low nuclear recoil (NR) energies [8]. In this Letter, the first detection of CENS induced by solar neutrinos with the XENONnT experiment is reported. This is a “first” in three different aspects: the first detection of elastic NRs from astrophysical neutrinos, the first measurement of the CENS process with a Xe target, and the first step into the “neutrino fog” by a DM experiment [9, 10].
@@
The XENONnT experiment [11], located at the INFN Laboratori Nazionali del Gran Sasso in Italy, is designed to search for Weakly Interacting Massive Particles (WIMPs) scattering off Xe nuclei, which has a similar NR signature as CENS. The experiment consists of three nested detectors: a muon veto (MV), a neutron veto (NV), and an innermost LXe detector. The latter is a two-phase time projection chamber (TPC) housed in a double-walled cryostat filled with 8.5 t of LXe. The cylindrical TPC, 1.33 m in diameter and 1.49 m in height, is enclosed by polytetrafluoroethylene (PTFE) panels and viewed by 494 3-inch Hamamatsu R11410-21 photomultiplier tubes (PMTs) [12] arranged in a top and a bottom array. The active LXe mass in the TPC is 5.9 t.
Particle interactions in the TPC produce both scintillation photons and ionization electrons. The prompt scintillation photons are detected by the PMTs and are referred to as the S1 signal. The liberated electrons drift upwards in the drift field to the liquid-gas interface, where they are extracted into the gas and produce a secondary scintillation signal, called the S2 signal, via electroluminescence. The time difference between and signals is proportional to the interaction depth (). Event positions in the horizontal plane () are reconstructed based on the hit patterns of S2 signals in the top PMT array.
@@
Dataset — This search uses two datasets with a total live time of 316.5 days after accounting for dead time from DAQ [14] and vetoes. The first dataset, taken between July 6, 2021, and November 10, 2021 (the first science run) [15] and in the following two weeks [15], is referred to as the SR0 dataset in this Letter, with a combined live time of 108.0 days. The second dataset is collected between May 19, 2022, and August 8, 2023, a period referred to as SR1, with a live time of 208.5 days. During SR0 (SR1), the temperature and pressure in the detector are stable within (176.80.4) K ((177.20.4) K) and (1.8900.004) bar ((1.920.02) bar), respectively. The liquid level in SR0 is stable within (5.020.20) mm [16]. On July 15, 2022, the liquid level is lowered by 0.2 mm and the anode voltage is raised by 50 V to mitigate localized electron bursts and maintain a consistent extraction field strength. Before and after this adjustment, the liquid level in SR1 is maintained stable at 5.0 mm and 4.8 mm above the gate electrode, respectively.
@@
Besides the 17 PMTs already excluded from the analyses during SR0 [17], three additional PMTs are removed in SR1 due to increased afterpulsing or intermittent light emission. PMT gains are monitored weekly using pulsed LED signals, and are found to be stable in SR0 (SR1) within 3% (3.5%). PMT hits are recorded on a per-PMT basis when crossing the digitization threshold, typically about 2.06 mV [14]. The mean single photoelectron (PE) acceptance in SR0 (SR1) is determined to be (91.20.2%) (). Clusters of PMT hits in time are divided into peaks, which are classified into S1 and S2 signals based on their waveforms and intensity distributions on PMT arrays [18, 19].
A distortion of the drift field near the edges of the detector leads to a difference in positions between the interaction site and the extraction position. It also leads to a small charge-insensitive volume (CIV) [13] in the lower part of the TPC, from where the drifting electrons reach the PTFE wall instead of the liquid-gas interface. A data-driven correction for the radial coordinate is applied to reproduce the uniform distribution of calibration events [20]. For SR0, the method from [15] is kept, where the CIV does not enter the position correction but is considered in the fiducial volume (FV) calculation. The FV mass uncertainty originated from field distortion and position reconstruction is less than 5%. In SR1, the event positions are corrected according to the boundary defined by the simulated drift field [13] to account for the CIV. After considering the field distortion correction and removing events with the interaction depth below -142 cm or above -13 cm due to an insufficient understanding of the detector and backgrounds, the FV mass for SR0 and SR1 are (3.97 0.20) t and (4.10 0.19) t, respectively. The total exposure in this analysis is 3.51 ty.
Light from S1 or S2 signals can create delayed electron signals via photoionization of impurities in the LXe [16]. The photoionization strength, defined as the ratio between the number of measured photoionization electrons within after an signal larger than and the number of electrons in the signal itself, increased tenfold after a long maintenance and upgrade phase between SR0 and SR1, including changing the radon removal system [21] to extract not only xenon from the gaseous but also from the liquid volume with high flow rate. One hypothesis is the release of photoionizable impurities from components of the system. No impact is observed from these impurities on the electron lifetime, which is an attenuation coefficient for the attachment to electronegative impurities during the drift of ionization electrons.
@@
After all corrections, the stability of the corrected and corrected (c) signals in SR0 are within 1% and 1.9%, respectively, and 0.3% and 1.1% in SR1. The variations are propagated as uncertainties into the determination of the photon gain (1) and electron gain (2). Using the method described in [15], 1 and 2 in SR0 (SR1) are found to be PE/photon ( PE/photon) and PE/electron ( PE/electron), respectively.
@@
The expected NR spectrum of CENS in LXe, considering the solar neutrino flux measured by the SNO experiment [22] and the CENS cross-section on Xe predicted by the SM [23], is shown in Fig. 1, with 90% of detectable recoils between 0.7 and 2.1 keV. The main contribution is from neutrinos with energies between 8 and 15 MeV. The low-energy NR response in this search is calibrated with 152 keV neutrons from an external source [24], with the recoil spectrum also shown in Fig. 1. The uncertainty in signal acceptance arises from uncertainties in S1 reconstruction, classification acceptance, and event selection acceptance. A model for light yield () and charge yield () is fitted [25] to calibration data using a method similar to that described in [26]. The uncertainties of yields are propagated into the final inference with two parameters, and . This calibration will be presented in an upcoming publication [27]. and below are assumed to be zero, which has a negligible impact on the CENS detection rate.
@@
The expected CENS rate in our previous WIMP search region [28] is 0.2 events/(ty). To increase the rate of detected CENS in this search, the signal acceptance is improved by lowering two thresholds. First, the S2 signal threshold is reduced from 200 PE in the WIMP search to 120 PE in this search. Second, the requirement of S1 coincidence, defined as the number of PMTs with hits within around the maximal amplitude of the S1 waveform, is lowered from three to two. The reduced thresholds lead to an expected CENS detection rate of 3.7(3.3) events/(ty) in SR0(1), a factor of 17 larger than in the WIMP search.
The ROI in this analysis is defined to be two or three hits for signals and (120, 500) PE for signals. The upper bound of the area range is set to retain most of CENS signal and to remove electronic recoil (ER) backgrounds from and radiation, which have higher ratios of S2 to S1 than NRs [29]. S1 signals with more than three hits are rarely produced by CENS and such events are therefore not included in this analysis. Events in the ROI are blinded except those with radial positions larger than 63.0 cm, which are used to model the surface events produced by 210Pb plate-out on the TPC wall [30], and are not part of the dataset for the search. Threefold events were unblinded in the SR0 WIMP search [13], which contributes to % of total CENS rate since twofold events dominates and SR1 has more exposure.
Cuts based on the features of S1 and S2 peaks, inherited from [17], are employed to ensure the quality of the reconstructed events. signals composed of at least two hits are required to be larger than 1 PE. signals up to 4 PE are accepted in size per PMT. S2 signals must be detected by both PMT arrays with a reasonable signal fraction in the top array. S2 signals detected on the top array are also required to follow the expected pattern from the optical response of XENONnT. Events with multiple S2 signals are rejected to suppress the neutron background. As in [17], events found in coincidence with either MV or NV are rejected.
@@
The AC is the dominant background, formed by accidentally paired “isolated” S1 and S2 signals. The accidental pile-up rate of these isolated S1 and S2 signals within the maximum drift time is significant, reaching several hundred events per day with no mitigation measures applied.
The primary source of the isolated S1 and S2 signals in the CENS search ROI are delayed signals after high-energy (HE) interactions. Those interactions, with characteristic S2 areas larger than induced predominantly by -rays, are known to contaminate their subsequent time interval with single photoelectron PMT hits and small S2 signals.
@@
The impact on an isolated signal by a preceding HE event is quantified by the ratio of to , where is the S2 area of the HE event and is the time between the HE event and the isolated signal. All the HE events one second before the isolated signal are considered and the event with the largest ratio of to (defined as ) is identified as the most influential one on the isolated signal rate. Cuts are then applied on to minimize the isolated signal rate. A time window of (one maximum drift time) is vetoed after any HE interaction in SR0. In SR1, due to the increased photoionization rate, the veto window is extended to . The cut on for 2 (3)-hit S1 signals is less than () PE/µs, effectively reducing isolated S1 rates by more than 80% (50%) while accepting 87% (96%) of CENS signals.
@@
For isolated S2 signals, correlations with preceding HE events and the localized SE burst in ()-position are utilized, accounting for the uncertainty in position reconstruction.
@@
By injecting simulated CENS signals at random times and positions into the real data, the overall acceptance of these cuts is evaluated to be 75% (85%) for 2 (3)-hit signals. The isolated S1 and S2 waveforms are then sampled and assigned a random drift time before being merged into artificial AC events. Facilitated by [34], the simulation improved compared to [6] in preserving the spectrum and modeling the time dependence to minimize the systematic uncertainties of the AC model.
@@
Consequently, events close to the perpendicular wires are excluded from the analysis. Due to the S2-area-dependent position resolution, this leads to an S2-area-dependent reduction in the S2 acceptance rather than a reduction of the fiducial mass.
@@
Surface events produced mainly by 210Pb plate-out on the TPC wall have reduced S2 signals [30], which could lead to leakage of events into the ROI. A data-driven approach is adopted to derive the radial distribution of this background. Due to the limited statistical data, deriving and validating the data-driven model across all four analysis dimensions is currently unfeasible. Consequently, the outer radius of the FV for SR0 (SR1) is set at (), such that surface events are expected to be less than 0.12 (0.23), respectively. At this level, this background can be safely neglected without risk of signal-like mis-modeling in the CENS search according to a dedicated toy Monte-Carlo (MC) study.
Radiogenic neutrons originating from the detector materials are modeled using the framework of [26] with neutron spectra from updated knowledge of the detector material radioactivity. The prediction for SR0 and SR1 are () events and () events, respectively. The rate uncertainty of 58% is derived from neutron candidates in SR0 tagged by the NV. A dedicated NV and MV tagged unblinding of the CEvNS ROI observed in total a single event per veto. This is in agreement with the expected number of events vetoed by accidental coincidence between the TPC and the veto detectors.
The ER background is composed mainly of -decays from radioactive impurities such as 214Pb and 85Kr, and electrons scattered by external gamma rays and solar neutrinos [26]. The shape of the ER background in the CENS ROI is generated by [25] with emission model fit to the 220Rn calibration data [26]. The rate of ER background events is derived by fitting the events with ER energy above 20 keV, assuming a flat ER spectrum. However, the emission model in low energy has large systematical uncertainty. If using the light and charge yields from NEST [36, 37], the expected ER rate is 10 times lower. To account for this discrepancy, a 100% uncertainty is assigned to the ER rate. Consequently, the assumed ER background in SR0 and SR1 is taken to be at most and events, respectively. Measurement of the light and charge response in XENONnT with a 0.27 keV calibration using a 37Ar electron capture (EC) source, which will be introduced in a future publication, also confirms the nominal value is a conservative choice.
@@
, BDT score, BDT score, and cS2 are the four parameters used to discriminate between the CENS signal and the dominating AC background. The background and signal models are coarsely binned, with three bins in each of the four analysis parameters for a total of 81 bins. A 4-dimensional binned likelihood analysis is performed. The bins are chosen to have the same expected number of AC background events in the projection of each parameter. The chance for mis-modeling of the AC background due to the limited number of isolated S1 and S2 peaks is negligible, as validated via toy-MC simulations.
@@
where the parameter of interest can either be the solar neutrino flux (), or the flux-weighted CENS cross-section on Xe (). are the nuisance parameters, iterates through the two science runs, and iterates through the nuisance terms: the constraints on and , the signal acceptance uncertainty and the uncertainties in the rates of the AC, neutron, and ER backgrounds. The nuisance parameters are constrained via external measurements, modeled by Gaussian pull terms . The models of CENS and neutrons change in shape and expectation value with and . The AC background rates are independent between science runs, while all other parameters are coupled.
The CENS discovery significance and the construction of a confidence interval for the neutrino flux are computed using a test statistic based on the profile log-likelihood ratio as in [38, 26]. The critical region for the confidence interval construction and expected discovery significance are computed with toy-MC simulations using [39]. Consistency between the model and data is evaluated by a combination of four binned likelihood GOF tests performed on the four one-dimensional projections, combining SR0 and SR1. The p-values are computed based on the distribution of the binned likelihood GOF test statistic obtained via toy-MC simulations. A threshold of 0.013 is selected for each test to obtain a 95% confidence limit (CL) for the final combined test. The test is defined before unblinding and its suitability to reject mis-modeling is assessed using toy-MC simulations.
The strategy to report the result from the CENS search is decided before unblinding. A Feldman-Cousins construction [40] is used to constrain the solar neutrino flux and the CENS cross-section without setting a threshold on p-values for reporting a two-sided measurement.
The expected CENS signal under the nominal emission model is events, with the uncertainty originating from S1 and S2 acceptances and detector response to low-energy NRs. The expected background is events, dominated by the AC background, as shown in Tab. 1. With the final background prediction summarized in Tab. 1, the probability of obtaining a (3) discovery significance with this dataset is estimated to be 80% (48%) using toy-MC simulations.
Figure 2: Distributions of best-fit signal and background, together with the data in the projected analysis dimensions, summing both science runs. The observed number of events with Poisson uncertainties in each bin is shown in black. The CENS signal is represented by the light green histogram on top of the backgrounds, which are indicated by purple (AC), blue (ER), and yellow (Neutron) histograms. As the bin edges on each analysis dimension vary from SR0 to SR1, the plot for c is shown in double-axis, and the other dimensions are shown in quantiles of the AC background for the summed results.
@@
The distributions of the observed 37 events and the best-fit model projected to each analysis dimension are shown in Fig. 2. The p-values in c, BDT score, and BDT score show a good match between the unconstrained best-fit model and observations. The p-value in the is 0.008, indicating a potential mis-modeling.
@@
Abandoning in the statistical inference would lead to a larger best-fit CENS signal of 13.1 events with a statistical significance of .
@@
Assuming the flux-weighted CENS cross-section predicted by the SM, Fig. 3 shows the XENONnT constraint on the solar neutrino flux of at 68% CL.

### v2

# First Indication of Solar 8B Neutrinos via Coherent Elastic Neutrino-Nucleus Scattering with XENONnT

@@
First Indication of Solar 8B Neutrinos via Coherent Elastic Neutrino-Nucleus Scattering with XENONnT
@@ Abstract
We present the first measurement of nuclear recoils from solar 8B neutrinos via coherent elastic neutrino-nucleus scattering with the XENONnT dark matter experiment. The central detector of XENONnT is a low-background, two-phase time projection chamber with a 5.9 t sensitive liquid xenon target. A blind analysis with an exposure of 3.51 tyr resulted in 37 observed events above 0.5 keV, with () events expected from backgrounds. The background-only hypothesis is rejected with a statistical significance of 2.73 . The measured 8B solar neutrino flux of is consistent with results from the Sudbury Neutrino Observatory. The measured neutrino flux-weighted CENS cross section on Xe of is consistent with the Standard Model prediction. This is the first direct measurement of nuclear recoils from solar neutrinos with a dark matter detector.
@@
Coherent elastic neutrino-nucleus scattering (CENS) is a Standard Model (SM) process with low momentum transfer, which allows neutrinos to scatter coherently with nuclei [1, 2, 3]. This process has only recently been observed using an intense, pulsed spallation neutron source (SNS) [4, 5]. The detection of CENS events from solar neutrinos is more challenging due to the lower flux [6] and energy, as well as the lack of timing information. Therefore, it requires minimal backgrounds and maximizing the sensitive region of interest (ROI) with a low-energy threshold. Liquid xenon (LXe) detectors searching for dark matter (DM) particles fulfill these requirements, but have not been able to reach the required sensitivity until now [7, 8]. Solar 8B neutrinos are expected to contribute the largest detectable number of coherent neutrino-xenon scattering events, albeit at low nuclear recoil (NR) energies [9]. In this Letter, the first detection of CENS induced by solar neutrinos with the XENONnT experiment is reported. This is a “first” in three different aspects: the first detection of elastic NRs from astrophysical neutrinos, the first measurement of the CENS process with a Xe target, and the first step into the “neutrino fog” by a DM experiment [10, 11].
@@
The XENONnT experiment [12], located at the INFN Laboratori Nazionali del Gran Sasso in Italy, is designed to search for weakly interacting massive particles (WIMPs) scattering off Xe nuclei, which has a similar NR signature as CENS. The experiment consists of three nested detectors: a muon veto (MV), a neutron veto (NV), and an innermost LXe detector. The latter is a two-phase time projection chamber (TPC) housed in a double-walled cryostat filled with 8.5 t of LXe. The cylindrical TPC, 1.33 m in diameter and 1.49 m in height, is enclosed by polytetrafluoroethylene (PTFE) panels and viewed by 494 3-in. Hamamatsu R11410-21 photomultiplier tubes (PMTs) [13] arranged in a top and a bottom array. The active LXe mass in the TPC is 5.9 t.
Particle interactions in the TPC produce both scintillation photons and ionization electrons. The prompt scintillation photons are detected by the PMTs and are referred to as the S1 signal. The liberated electrons drift upward in the drift field to the liquid-gas interface, where they are extracted into the gas and produce a secondary scintillation signal, called the S2 signal, via electroluminescence. The time difference between and signals is proportional to the interaction depth (). Event positions in the horizontal plane () are reconstructed based on the hit patterns of S2 signals in the top PMT array.
@@
Dataset — This search uses two datasets with a total live time of 316.5 days after accounting for dead time from data acquisition [15] and vetoes. The first dataset, taken between July 6, 2021 and November 28, 2021 is referred to as the SR0 dataset in this Letter with a live time of 108.0 days. The second dataset was collected between May 19, 2022 and August 8, 2023, a period referred to as SR1, with a live time of 208.5 days. During SR0 (SR1), the temperature and pressure in the detector are stable within (176.80.4) [(177.20.4)] K and (1.8900.004) [(1.920.02)] bar, respectively. The liquid level in SR0 is stable within (5.020.20) mm [16]. On July 15, 2022, the liquid level is lowered by 0.2 mm and the anode voltage is raised by 50 V to mitigate localized electron bursts and maintain a consistent extraction field strength. Before and after this adjustment, the liquid level in SR1 is maintained stable at 5.0 and 4.8 mm above the gate electrode, respectively.
@@
In addition to the 17 PMTs already excluded from the analyses during SR0 [17], three additional PMTs are removed in SR1 due to increased afterpulsing or intermittent light emission. PMT gains are monitored weekly using pulsed LED signals, and are found to be stable in SR0 (SR1) within 3% (3.5%). PMT hits are recorded on a per-PMT basis when crossing the digitization threshold, typically about 2.06 mV [15]. The mean single photoelectron (PE) acceptance in SR0 (SR1) is determined to be (91.20.2%) [(92.10.7) %]. Clusters of PMT hits in time are divided into peaks, which are classified into S1 and S2 signals based on their waveforms and intensity distributions on PMT arrays [18, 19].
A distortion of the drift field near the edges of the detector leads to a difference in positions between the interaction site and the extraction position. It also leads to a small charge-insensitive volume (CIV) [14] in the lower part of the TPC, from where the drifting electrons reach the PTFE wall instead of the liquid-gas interface. A data-driven correction for the radial coordinate is applied to reproduce the uniform distribution of calibration events [20]. For SR0, the method from [21] is kept, where the CIV does not enter the position correction but is considered in the fiducial volume (FV) calculation. The FV mass uncertainty originated from field distortion and position reconstruction is less than 5%. In SR1, the event positions are corrected according to the boundary defined by the simulated drift field [14] to account for the CIV. After considering the field distortion correction and removing events with the interaction depth below -142 cm or above -13 cm due to an insufficient understanding of the detector and backgrounds, the FV mass for SR0 and SR1 are (3.97 0.20) and (4.10 0.19) t, respectively. The total exposure in this analysis is 3.51 tyr.
Light from S1 or S2 signals can create delayed electron signals via photoionization of impurities in the LXe [16]. The photoionization strength, defined as the ratio between the number of measured photoionization electrons within after an signal larger than and the number of electrons in the signal itself, increased tenfold after a long maintenance and upgrade phase between SR0 and SR1. One hypothesis of the increased photoionization is that components in the radon removal system [22] are releasing photoionizable impurities after the upgrade, which enabled high flow extraction from the LXe target. No impact is observed from these impurities on the electron lifetime, which is an attenuation coefficient for the attachment to electronegative impurities during the drift of ionization electrons.
@@
After all corrections, the stability of the corrected and corrected (c) signals in SR0 are within 1% and 1.9%, respectively, and 0.3% and 1.1% in SR1. The variations are propagated as uncertainties into the determination of the photon gain (1) and electron gain (2). Using the method described in [21], 1 and 2 in SR0 (SR1) are found to be (0.1510.001) [(0.1370.001)] PE/photon and (16.50.6) [(16.90.5)] PE/electron, respectively.
@@
The expected NR spectrum of CENS in LXe, considering the solar neutrino flux measured by the Sudbury Neutrino Observatory(SNO) [23], the neutrino energy spectrum from [24] and the CENS cross section on Xe predicted by the SM [25], is shown in Fig. 1, with 90% of detectable recoils between 0.7 and 2.1 keV. The main contribution is from neutrinos with energies between 8 and 15 MeV. The low-energy NR response in this search is calibrated with 152 keV neutrons from an external source [26], with the recoil spectrum also shown in Fig. 1. The uncertainty in signal acceptance arises from uncertainties in S1 reconstruction, classification acceptance, and event selection acceptance. A model for light yield () and charge yield () is fitted [27] to calibration data using a method similar to that described in [28]. The uncertainties of yields are propagated into the final inference with two parameters, and , which determine the relative shift of and from their median toward the 1 quantiles. This calibration will be presented in an upcoming publication [29]. and below are assumed to be zero, which has a negligible impact on the CENS detection rate.
@@
The expected CENS rate in our previous WIMP search region [30] is 0.2 events/(tyr). To increase the rate of detected CENS in this search, the signal acceptance is improved by lowering two thresholds. First, the S2 signal threshold is reduced from 200 PE in the WIMP search to 120 PE in this search. Second, the S1 coincidence requirement was lowered from threefold coincidence to twofold coincidence, now minimally requiring only two PMTs with hits within around the maximal amplitude of the S1 waveform. The reduced thresholds lead to an expected CENS detection rate of 3.7(3.3) events/(tyr) in SR0(1), a factor of 17 larger than in the WIMP search.
The ROI in this analysis is defined to be two or three hits for signals and (120, 500) PE for signals. The upper bound of the area range is set to retain most of CENS signal and to remove electronic recoil (ER) backgrounds from and radiation, which have higher ratios of S2 to S1 than NRs [31]. S1 signals with more than three hits are rarely produced by CENS and such events are therefore not included in this analysis. Events in the ROI are blinded except those with radial positions larger than 63.0 cm, which are used to model the surface events produced by 210Pb plate out on the TPC wall [32], and are not part of the dataset for the search. Threefold events were unblinded in the SR0 WIMP search [14], which contributes to % of total CENS rate since twofold events dominate and SR1 has more exposure.
Cuts based on the features of S1 and S2 peaks, inherited from [17], are employed to ensure the quality of the reconstructed events. signals composed of at least two hits are required to be larger than 1 PE. signals up to 4 PE are accepted in size per PMT. S2 signals must be detected by both PMT arrays with a reasonable signal fraction of around 75% in the top array. S2 signals detected on the top array are also required to follow the expected pattern from the optical response of XENONnT. Events with multiple S2 signals are rejected to suppress the neutron background. As in [17], events found in coincidence with either MV or NV are rejected.
@@
The AC is the dominant background, formed by accidentally paired “isolated” S1 and S2 signals. The accidental pileup rate of these isolated S1 and S2 signals within the maximum drift time is significant, reaching several hundred events per day before mitigation measures are applied.
The primary source of the isolated S1 and S2 signals in the CENS search ROI are delayed signals after high-energy (HE) interactions. These interactions, with characteristic S2 areas larger than induced predominantly by rays from the materials’ radioactivity, are known to contaminate their subsequent time interval with single photoelectron PMT hits and small S2 signals.
@@
The impact on an isolated signal by a preceding HE event is quantified by the ratio of to , where is the S2 area of the HE event and is the time between the HE event and the isolated signal. All the HE events 1 sec before the isolated signal are considered and the event with the largest ratio of to (defined as ) is identified as the most influential one on the isolated signal rate. Cuts are then applied on to minimize the isolated signal rate. A time window of (one maximum drift time) is vetoed after any HE interaction in SR0. In SR1, due to the increased photoionization rate, the veto window is extended to . The cut on for 2 (3)-hit S1 signals is less than () PE/µs, effectively reducing isolated S1 rates by more than 80% (50%) while accepting 87% (96%) of CENS signals.
@@
For isolated S2 signals, correlations with preceding HE events and the localized SE burst in () position are utilized, accounting for the uncertainty in position reconstruction.
@@
By injecting simulated CENS signals at random times and positions into the real data, the overall acceptance of these cuts is evaluated to be 75% (85%) for 2- (3)-hit signals. The isolated S1 and S2 waveforms are then sampled and assigned a random drift time before being merged into artificial AC events. Facilitated by [36], the simulation improved compared to [7] in preserving the spectrum and modeling the time dependence to minimize the systematic uncertainties of the AC model.
@@
Consequently, events close to the perpendicular wires are excluded from the analysis. Because of the S2-area-dependent position resolution, this leads to an S2-area-dependent reduction in the S2 acceptance rather than a reduction of the fiducial mass.
@@
Surface events produced mainly by 210Pb plate out on the TPC wall have reduced S2 signals [32], which could lead to leakage of events into the ROI. A data-driven approach is adopted to derive the radial distribution of this background. Because of the limited statistical data, deriving and validating the data-driven model across all four analysis dimensions is currently unfeasible. Consequently, the outer radius of the FV for SR0 (SR1) is set at (), such that surface events are expected to be less than 0.12 (0.23), respectively. At this level, this background can be safely neglected without risk of signal-like mis-modeling in the CENS search according to a dedicated toy Monte Carlo (MC) study.
Radiogenic neutrons originating from the detector materials are modeled using the framework of [28] with neutron spectra from updated knowledge of the detector material radioactivity. The prediction for SR0 and SR1 are () and () events, respectively. The rate uncertainty of 58% is derived from neutron candidates in SR0 tagged by the NV. In the CEvNS ROI, the NV and MV tagged one event each after a dedicated unblinding, which is in agreement with the expected number of events vetoed by accidental coincidence between the TPC and the veto detectors.
The ER background is composed mainly of decays from radioactive impurities such as 214Pb and 85Kr, and electrons scattered by external rays and solar neutrinos [28]. The shape of the ER background in the CENS ROI is generated by [27] with emission model fit to the 220Rn calibration data [28]. The rate of ER background events is derived by fitting the events with ER energy above 20 keV, assuming a flat ER spectrum. However, the emission model in low energy has large systematical uncertainty. If using the light and charge yields from the Noble Element Simulation Technique(NEST) [38, 39], the expected ER rate is 10 times lower. To account for this discrepancy, a 100% uncertainty is assigned to the ER rate. Consequently, the assumed ER background in SR0 and SR1 is taken to be at most and events, respectively. Measurement of the light and charge response in XENONnT with a 0.27 keV calibration using a 37Ar electron capture (EC) source, which will be introduced in a future publication, also confirms the nominal rate of the ER background is a conservative choice.
@@
, BDT score, BDT score, and cS2 are the four dimensions used to discriminate between the CENS signal and the dominating AC background. The background and signal models are coarsely binned, with three bins in each of the four analysis dimensions for a total of 81 bins. A four-dimensional binned likelihood analysis is performed. The bins are chosen to have the same expected number of AC background events in the projection of each dimension. The chance for mis-modeling of the AC background due to the limited number of isolated S1 and S2 peaks is negligible, as validated via toy MC simulations.
@@
where the parameter of interest can either be the solar neutrino flux (), or the flux-weighted CENS cross section on Xe (). are the nuisance parameters, iterates through the two science runs, and iterates through the nuisance terms: the constraints on and , the signal acceptance uncertainty and the uncertainties in the rates of the AC, neutron, and ER backgrounds. The nuisance parameters are constrained via external measurements, modeled by Gaussian pull terms . The models of CENS and neutrons change in shape and expectation value with and . The AC background rates are independent between science runs, while all other parameters are coupled.
The CENS discovery significance and the construction of a confidence interval for the neutrino flux are computed using a test statistic based on the profile log-likelihood ratio as in [40, 28]. The critical region for the confidence interval construction and expected discovery significance are computed with toy MC simulations using [41]. Consistency between the model and data is evaluated by a combination of four binned likelihood GOF tests performed on the four one-dimensional projections, combining SR0 and SR1. The p-values are computed based on the distribution of the binned likelihood GOF test statistic obtained via toy MC simulations. A threshold of 0.013 is selected for each test to obtain a 95% confidence limit (CL) for the final combined test. The test is defined before unblinding and its suitability to reject mis-modeling is assessed using toy MC simulations.
The strategy to report the result from the CENS search is decided before unblinding. A Feldman-Cousins construction [42] is used to constrain the solar neutrino flux and the CENS cross section without setting a threshold on p-values for reporting a two-sided measurement.
The expected CENS signal under the nominal emission model is events, with the uncertainty originating from S1 and S2 acceptances and detector response to low-energy NRs. The expected background is events, dominated by the AC background, as shown in Tab. 1. With the final background prediction summarized in Tab. 1, the probability of obtaining a (3) discovery significance with this dataset is estimated to be 80% (48%) using toy MC simulations.
Figure 2: Distributions of best-fit signal and background, together with the data in the projected analysis dimensions, summing both science runs. The observed number of events with Poisson uncertainties in each bin is shown in black. The CENS signal is represented by the light green histogram on top of the backgrounds, which are indicated by purple (AC), blue (ER), and yellow (neutron) histograms. As the bin edges on each analysis dimension vary from SR0 to SR1, the plot for c is shown in double axis, and the other dimensions are shown in quantiles of the AC background for the summed results.
@@
The distributions of the observed 37 events and the best-fit model projected to each analysis dimension are shown in Fig. 2. A detailed plot showing the SR0 and SR1 results separately is presented in Appendix 6. The p-values in c, BDT score, and BDT score show a good match between the unconstrained best-fit model and observations. The p-value in the is 0.008, indicating a potential mis-modeling.
@@
Abandoning in the statistical inference would lead to a larger best-fit CENS signal of 13.1 events with a statistical significance of . In addition, two tests of overdensity in () space were defined before unblinding, although not part of the analysis dimensions. One returned a p-value below the threshold of 0.018, prompting checks including inspection of event distributions in all cut spaces that show no indication of mis-modeling.
@@
Assuming the flux-weighted CENS cross section predicted by the SM, Fig. 3 shows the XENONnT constraint on the solar neutrino flux of at 68% CL.